I am having troubles with this one, could you please help?
What is the y and x-intercept of the line 2y=3x-6?
Thanks,
Lackinmath

Answer by checkley71(8403)

Y=mX+b
2Y=3X-6
Y=3X/2-6/2
Y=3X/2-3 SO THE Y INTERCEPT IS (0,-3)
TO FIND THE X INTERCEPT SET Y=0 AND SOLVE
0=3X/2-3
3X/2=3
3X=6
X=6/3
X=2 THIS IS THE X INTERCEPT(2,0)
###########################################################
MERRY CHRISTMAS.